A researcher conducts a mileage economy test Involving 107 cars. The frequency distribution describing average mlles per gallon mpg) appears In the following table. polnts 15 up to 28 28 up to 25 25 up to 38 38 up to 35 З5 up to 4e 48 up to 45 15 26 24 13 eBook Refeencea. Construct the corresponding relative frequency, cumulative frequency, and cumulative relative frequency distributions. (Round Relative Frequency and Cumulative Relative Frequency to 4 decimal places.) Relative Cumulative Relative Average mpg 15 up to 20 20 up to 25 25 up to 30 30 up to 35 35 up to 40 40 up to 45 b-1. How many of the cars got less than 20 mpg? Number of cars b-2 What percentage of the cars got at least 25 but less than 30 mpg? (Round your answer to 2 decimal places.) Percentage of cars b-3. What percentage of the cars got less than 30 mpg? (Round уоиг answer to 2 decimal places.) Percentage of cars b-4. What percent got 30 mpg or more? (Round your answer to 2 decimal places.) Percentage of cars
Answer #1

Given that,

a )

Class Frequency Relative Frequency Cumulative Frequency Cumulative Relative Frequency
15-20 15 15 / 107 = 0.1402 15 15 / 107 = 0.1402
20-25 26 26 / 107= 0.2430 15+26 = 41 41 / 107 = 0.3832
25-30 24 24 /107 = 0.2243 41 +24 = 65 65 / 107 = 0.6075
30-35 13 13 / 107 = 0.1215 65 +13 = 78 78 /107 = 0.7290
35-40 7 7 / 107 = 0.0654 78 + 7 = 85 85 /107 = 0.7944
40-45 22 22 /107= 0.2056 85 + 22 = 107 107 / 107 = 1.0000
Total 107 1

b 1 ) The cars got less than 20 mpg are 15 cars

b 2 ) The cars got at least 25 mpg but less than 30 mpg are 22.43% cars

b 3 ) The cars got less than 30 mpg are 60.75% cars

b 4 ) The cars got or more than 30 mpg are 60.75% cars